OpenStudy (anonymous):

(1/2^1/2)^2= that is square root of one over 2 square is (1/2^1/2)^2= that is square root of one over 2 square is @Mathematics

OpenStudy (anonymous):

\[\frac{1}{2}\]

OpenStudy (anonymous):

but it should be 1/4

OpenStudy (agreene):

\[\frac{\sqrt{1}}{2^2}\] Is that your question? what you said and what you typed are rather different.

OpenStudy (asnaseer):

is this your question - what is the square root of:\[\frac{1}{2^2}\]

OpenStudy (asnaseer):

or is it this - what is the square root of:\[(\frac{1}{\sqrt{2}})^2\]

OpenStudy (anonymous):

yeah asnaseer the 2nd one

OpenStudy (agreene):

If it should be 1/4 I assume its: \[\frac{\sqrt{1}}{2^2}=\frac{1}{4}\] sqrt 1 = 1 2^2 = 4 Of course, it could be: \[(\frac{\sqrt 1}{2})^2=(\frac{1}{2})^2=\frac{1}{4}\]

OpenStudy (agreene):

(1/(sqrt 2))^2 = 1/2 which fails the previous condition you gave that it needs to be 1/4 and sqrt (1/(sqrt 2))^2 = 1/(sqrt 2)

OpenStudy (asnaseer):

in that case just remember that square root is the inverse of square (with a plus/minus of course), i.e.:\[\sqrt{x^2}=\pm x\]so:\[\sqrt{(\frac{1}{\sqrt{2}})^2}=\pm\frac{1}{\sqrt{2}}\]

OpenStudy (asnaseer):

because:\[x*x=x^2\]and:\[(-x)*(-x)=x^2\]

OpenStudy (anonymous):

i mean 1 divided by square root of all squared

OpenStudy (anonymous):

??

OpenStudy (anonymous):

i divided by square root of 2 all squared

OpenStudy (anonymous):

\[\sqrt{\frac{1}{2}}^2=\frac{1}{2}\]

OpenStudy (anonymous):

NO!!!!!!!!! 1 DIvided by square root of 2 then all squared

OpenStudy (anonymous):

\[(\frac{1}{\sqrt{2}})^2=\frac{1}{2}\] if you prefer

OpenStudy (anonymous):

same thing since \[\sqrt{\frac{1}{2}}=\frac{\sqrt{1}}{\sqrt{2}}=\frac{1}{\sqrt{2}}\]

OpenStudy (asnaseer):

ok, in that case:\[(\frac{1}{\sqrt{2}})^2=\frac{1}{\sqrt{2}}*\frac{1}{\sqrt{2}}=\frac{1*1}{\sqrt{2}*\sqrt{2}}=\frac{1}{2}\]

OpenStudy (anonymous):

yeah that one

OpenStudy (anonymous):

but why should it be 1/2?

OpenStudy (anonymous):

if the square roots cancel out,it should be 1/4

OpenStudy (asnaseer):

are you familiar with the rules for indexes? i.e. do you know:\[a^n*a^m=a^{n+m}\]

OpenStudy (anonymous):

yeah i do

OpenStudy (asnaseer):

if you do, then you could re-write:\[\sqrt{2}=2^{0.5}\]so that:\[\sqrt{2}*\sqrt{2}=2^{0.5}*2^{0.5}=2^{0.5+0.5}=2^1=2\]

OpenStudy (asnaseer):

@krypton - do you understand now?

OpenStudy (anonymous):

yeah i now do thanks

OpenStudy (asnaseer):

no problem - glad to help

OpenStudy (anonymous):

actually what you really need is that given any positive number, the square of the square root is that number. the square root means "the number whose square is given" so for example \[\sqrt{25}=5\iff 5^2=25\] therefore it is always true that for any positive number \[\sqrt{a}^2=a\]

OpenStudy (anonymous):

wolfram alpha people wolfram alpha